Question 1:

Multistage Growth At the beginning of 2010 stock of Boatright Stores is projecting the following dividends

Year Dividend
2010 0.84
2011 0.96
2012 1.16

After this dividend growth will be steady. Value Line forecasts a retention rate of 80% and a long term ROE of 12%. The required return for Boatright is 16%. The intrinsic value of Boatright is ________.

Question 2:

Free Cash Flow to Equity This year a firm has FCFF of $12.00 million. The firm has interest expense of $5 million and is in a 39% tax bracket and debt increased by $5 million. The firm's free cash flow to equity is ____________ million.

$12.00

$2.00

$8.95

$13.95
Answer #1

1). g = ROE x Retention Rate = 12% x 0.80 = 9.60%

P0 = [D1 / (1 + r)] + [D2 / (1 + r)2] + [D3 / (1 + r)3] + [{D3 x (1 + g)} / {(r - gC)(1 + r)3}]

= [$0.84 / 1.16] + [$0.96 / 1.162] + [$1.16 / 1.163] + [($1.16 x 1.096) / {(0.16 - 0.096) x 1.163}]

= $0.72 + $0.71 + $0.74 + $12.73 = $14.91

2). FCFE = FCFF – [Interest x (1-tax)] + Net Borrowings

= 12 - [5 x (1 - 0.39)] + 5 = 12 - 3.05 + 5 = $13.95 million

Hence, Option "D" is correct.